# Beyond Grids: Multi-objective Bayesian Optimization With Adaptive Discretization

### v3

@@ A.1 Preliminary results
Next, we state a proposition that relates the information gain with the posterior variance of the GP after each evaluation. This proposition will help us in obtaining information-type bounds on the sample complexity. Since its proof is lengthy, we defer it to Section A.9.
@@ A.2 Termination condition
(Termination condition for Adaptive $\bm{\epsilon}$-PAL) Let $\epsilon=\min_{j}\epsilon^{j}>0$, where $(\epsilon^{1},\ldots,\epsilon^{m})=\bm{\epsilon}$. When running Adaptive $\bm{\epsilon}$-PAL, if $\overline{\omega}_{t}\leq\epsilon$ holds at round $t$, then the algorithm terminates without further sampling.
@@
If $\overline{\omega}_{t}\leq\epsilon$ holds at iteration $t$, then for all $x_{h,i}\in{\cal S}_{t,0}\setminus{\cal P}_{t,2}$, we have $x_{h,i}\notin{\cal S}_{t,1}$.
@@
We have $\left\lVert\bm{\omega}_{h,i}\right\rVert_{2}\leq\overline{\omega}_{t}\leq\epsilon$, which implies that $\bm{\omega}_{h,i}\preceq\bm{\epsilon}$ since $\omega^{j}_{h,i}\leq\sqrt{\sum^{m}_{j^{\prime}=1}(\omega^{j^{\prime}}_{h,i})^{2}}\leq\epsilon\leq\epsilon^{j}$ for all $j\in[m]$.
@@
Since ${\cal S}_{t,1}\cup{\cal P}_{t,1}={\cal S}_{t,2}\cup{\cal P}_{t,2}$, we have for all $y\in{\cal S}_{t,1}\cup{\cal P}_{t,1}$
$\max(\bm{R}_{t}(y))-\min(\bm{R}_{t}(y))\preceq\bm{\epsilon}~{}.$ (4)
@@
$\max(\bm{R}_{t}(x_{h,i}))=\min(\bm{R}_{t}(x_{h,i}))+\bm{\omega}_{h,i}$ $\preceq\min(\bm{R}_{t}(x_{h,i}))+\bm{\epsilon}$
@@
$\min(\bm{R}_{t}(x_{h,i}))\preceq\min(\bm{R}_{t}(y^{*}))~{}.$ (6)
Since $y^{*}\in{\cal S}_{t,0}\cup{\cal P}_{t,0}$, by Definition 10 and equation 6, we conclude that $x_{h,i}\notin{\cal P}_{pess,t}$. Thus, we must have $x_{h,i}\in S_{t,0}\setminus{\cal P}_{pess,t}$. Finally, we claim that $\max(\bm{R}_{t}(x_{h,i}))\preceq_{\bm{\epsilon}}\min(\bm{R}_{t}(y^{**}))$ for some $y^{**}\in{\cal P}_{pess,t}$. Since equation 5 implies that $\max(\bm{R}_{t}(x_{h,i}))\preceq_{\bm{\epsilon}}\min(\bm{R}_{t}(y^{*}))$, if $y^{*}\in{\cal P}_{pess,t}$, then we can simply set $y^{**}=y^{*}$. Else if $y^{*}\notin{\cal P}_{pess,t}$, then this will imply by Definition 10 existence of $y^{**}\in{\cal P}_{pess,t}$ such that $\min(\bm{R}_{t}(y^{*}))\preceq\min(\bm{R}_{t}(y^{**}))$, which in turn together with equation 5 implies that $\max(\bm{R}_{t}(x_{h,i}))\preceq_{\bm{\epsilon}}\min(\bm{R}_{t}(y^{**}))$. Then, by the discarding rule of Adaptive $\bm{\epsilon}$-PAL specified in line 9 of Algorithm 1, we must have $x_{h,i}$ discarded by the end of the discarding phase, which implies that $x_{h,i}$ cannot be in $S_{t,1}$. This proves that all $x_{h,i}\in{\cal S}_{t,0}\setminus{\cal P}_{t,2}$ must be discarded.
@@
A.3 Guarantees on the termination of Adaptive $\bm{\epsilon}$-PAL
We start by stating a bound on the maximum depth node that can be created by Adaptive $\bm{\epsilon}$-PAL before it terminates for a given $\bm{\epsilon}$. This result will be used in defining “good" events that hold with high probability under which $\bm{f}(x_{h,i})$ lies in the confidence hyper-rectangle of $x_{h,i}$ formed by Adaptive $\bm{\epsilon}$-PAL, for all possible nodes that can be created by the algorithm (see Section A.4). Our bounds on sample complexity will hold given that these “good" events happen.
@@ A.3 Guarantees on the termination of Adaptive $\bm{\epsilon}$-PAL
Proof. By Lemma 3, at the latest, the algorithm terminates at round $t$ for which $\overline{\omega}_{t}\leq\epsilon$, i.e., $\overline{\omega}^{2}_{t}\leq\epsilon^{2}$. By definition, at a refining round we have
@@
$16mV^{2}_{h_{t}}\leq\epsilon^{2}~{},$
then we guarantee termination of the algorithm. We let
@@
for positive constants $C_{2}$ and $C_{3}$ defined in Corollary 1. Obviously, $V_{h}$ decays to 00 exponentially in $h$. Thus, by letting $h_{max}=h_{max}(\epsilon)$ to be the smallest $h\geq 0$, for which $16mV^{2}_{h_{max}}\leq\epsilon^{2}$ holds, it is observed that the algorithm stops refining at level $h_{max}$. ∎
@@
A.4 Two “good” events under which the sample complexity will be bounded
@@
(The first “good” event) For any $\delta\in(0,1)$, the probability of the following event is at least $21-\delta/$ 2:
@@
1($1-\mathbb{P}(\mathcal{F}_{1})$ $=\mathbb{E}\left[\mathbb{I}\left(\exists j\in[m],\exists\tau\geq 0,\exists x\in\mathcal{T}_{h_{max}}:|f^{j}(x)-\mu^{j}_{\tau}(x)|>\beta^{1/2}_{\tau}\sigma^{j}_{\tau}(x)\right)\right]$
@@
(The second “good” event) For any $\delta\in(0,1)$, the probability of the following event is at least $21-\delta/$ 2:
@@
$1-\mathbb{P}(\mathcal{F}_{2})$ =$=\mathbb{P}\left\{\exists h\geq 0,\exists i\in[N^{h}],\exists j\in[m]:\sup_{x\in B(x_{h,i},v_{1}\rho^{h},d)}|f^{j}(x)-f^{j}(x_{h,i})|>\omega(C_{\bm{k}}(v_{1}\rho^{h})^{\alpha})\right\}$
@@
A.5 Key results that hold under the “good” events
@@
In general, let the finite sequence of nodes $[y_{1},y_{2},\ldots,y_{n}]^{\mathsf{T}}$ be such that all the nodes $y_{1}$ to $y_{n-1}$ are discarded, meaning that, for each $i\in[n-1]$, the node $y_{i}$ stopped being part of $\mathcal{P}_{pess,s_{i+1}}$ at some round $s_{i+1}$ by satisfying
$\min(\bm{R}_{s_{i+1}}(c_{s_{i+1}}(y_{i})))\preceq\min(\bm{R}_{s_{i+1}}(y_{i+1}))~{}.$
@@
$\max(\bm{R}_{s_{1}}(c_{s_{1}}(x^{*})))-\bm{\epsilon}$ $\preceq\min(\bm{R}_{s_{1}}(y_{1}))\preceq\min(\bm{R}_{s_{2}}(c_{s_{2}}(y_{1})))$
$\preceq\min(\bm{R}_{s_{2}}(y_{2}))\preceq\min(\bm{R}_{s_{3}}(c_{s_{3}}(y_{2})))$
[…]
$\preceq\min(\bm{R}_{s_{n}}(y_{n}))\preceq\min(\bm{R}_{t}(c_{t}(y_{n})))~{},$
@@
$\max(\bm{R}_{s_{1}}(c_{s_{1}}(x^{*})))-\bm{\epsilon}\preceq\min(\bm{R}_{t}(c_{t}(y_{n})))\preceq\max(\bm{R}_{t}(c_{t}(y_{n})))~{}.$
@@
Suppose that $z_{2}\notin\hat{P}$. Then, the above process continues in a similar fashion. Suppose that the process never yields a node in $\hat{P}$. Since the algorithm is guaranteed to terminate by Lemma 3, the process stops and yields a finite sequence $[z_{1},\ldots,z_{n}]^{\mathsf{T}}$ of designs such that $z_{1}\notin\hat{P},\ldots,z_{n}\notin\hat{P}$. For each $i\in[n-1]$, let $s_{i+1}\in\mathbb{N}$ be the round at which $c_{s_{i+1}}(z_{i})$ is discarded; by the above process, we have
@@
$\min(\bm{R}_{s_{n+1}}(c_{s_{n+1}}(z_{n})))\preceq\min(\bm{R}_{s_{n+1}}(z_{n+1}))~{}.$ (20)
Hence, equation 20 and the earlier inequalities imply that
@@
where we have again used the shrinking property of the cumulative confidence hyper-rectangles in the second inequality. This means that $z_{n+1}$ is another node that $\bm{\epsilon}$-dominates $x$ but it is not in the finite sequence $[z_{1},\ldots,z_{n}]^{\mathsf{T}}$. This contradicts our assumption that the process stops after finding $n$ designs. Hence, at least one of the designs in $[z_{1},\ldots,z_{n}]^{\mathsf{T}}$ is in $\hat{P}$; let us call it $z$. This completes the proof of the claim.
@@
This shows that for every $\bm{\mu}\in\bm{f}(\mathcal{O}(\mathcal{X}))-\mathbb{R}^{m}_{+}$, there exists $z\in\hat{P}$ such that $\bm{\mu}\preceq_{\bm{\epsilon}}\bm{f}(z)$. By Definition 3, we have $\mathcal{Z}_{\bm{\epsilon}}(\mathcal{X})\subseteq\bm{f}(\mathcal{O}(\mathcal{X}))-\mathbb{R}^{m}_{+}$. Hence, for every $\bm{\mu}\in\mathcal{Z}_{\bm{\epsilon}}(\mathcal{X})$, there exists $\bm{\mu}^{\prime\prime}\in\bm{f}(\hat{P})$ such that $\bm{\mu}\preceq_{\bm{\epsilon}}\bm{\mu}^{\prime\prime}$. On the other hand, since we work under $\mathcal{F}_{1}\cap\mathcal{F}_{2}$, Lemma 9 implies that $\bm{f}(\hat{P})\subseteq\mathcal{Z}_{\bm{\epsilon}}(\mathcal{X})$. Therefore, $\bm{f}(\hat{P})$ is an $\bm{\epsilon}$-covering of $\mathcal{Z}_{\bm{\epsilon}}(\mathcal{X})$ (Definition 4 ), that is, $\hat{P}$ is an $\bm{\epsilon}$-accurate Pareto set (Definition 5). ∎
A.6 Derivation of the information-type sample complexity bound
@@
Let $\tau\in[\tau_{s}]$ and $j\in[m]$. Note that the $\tau$th evaluation is made at round $t_{\tau}$ and we have $\tau_{t_{\tau}}=\tau-1$. Hence, we have
@@
$\sqrt{\frac{16\beta_{T}\sigma^{2}Cm\gamma_{T}}{T}}\leq\epsilon~{}.$
@@
Proof. According to Lemma 3, we have $\overline{w}_{t_{s}}\leq\epsilon$. In addition, Lemma 13 says that
@@
$T=\min\left\{\tau\in\mathbb{N}:\sqrt{\frac{16\beta_{\tau}\sigma^{2}Cm\gamma_{\tau}}{\tau}}\leq\epsilon\right\}~{}.$
Since the event $\{\tau_{s}=T\}$ implies that $\{\overline{w}_{t_{s}}\leq\epsilon\}$, we have $\Pr({\tau_{s}}>T)=0$.∎
A.7 Derivation of the dimension-type sample complexity bound

### v4

@@ B.1 Preliminary results
Next, we state a proposition that relates the information gain with the posterior variance of the GP after each evaluation. This proposition will help us in obtaining information-type bounds on the sample complexity. Since its proof is lengthy, we defer it to Section B.9.
@@ B.2 Termination condition
(Termination condition for Adaptive $\bm{\epsilon}$-PAL) Let $\epsilon=\min_{j}\epsilon^{j}>0$, where $(\epsilon^{1},\ldots,\epsilon^{m})=\bm{\epsilon}$. When running Adaptive $\bm{\epsilon}$-PAL, if $\overline{\omega}_{t}<\epsilon$ holds at round $t$, then the algorithm terminates without further sampling.
@@
If $\overline{\omega}_{t}<\epsilon$ holds at iteration $t$, then for all $x_{h,i}\in{\cal S}_{t,0}\setminus{\cal P}_{t,2}$, we have $x_{h,i}\notin{\cal S}_{t,1}$.
@@
We have $\left\lVert\bm{\omega}_{h,i}\right\rVert_{2}\leq\overline{\omega}_{t}<\epsilon$, which implies that $\bm{\omega}_{h,i}\prec\bm{\epsilon}$ since $\omega^{j}_{h,i}\leq\sqrt{\sum^{m}_{j^{\prime}=1}(\omega^{j^{\prime}}_{h,i})^{2}}<\epsilon\leq\epsilon^{j}$ for all $j\in[m]$.
@@
Since ${\cal S}_{t,1}\cup{\cal P}_{t,1}={\cal S}_{t,2}\cup{\cal P}_{t,2}$, we have, for all $y\in{\cal S}_{t,1}\cup{\cal P}_{t,1}$, that
$\max(\bm{R}_{t}(y))-\min(\bm{R}_{t}(y))\prec\bm{\epsilon}~.$ (4)
@@
$\max(\bm{R}_{t}(x_{h,i}))=\min(\bm{R}_{t}(x_{h,i}))+\bm{\omega}_{h,i}$ $\prec\min(\bm{R}_{t}(x_{h,i}))+\bm{\epsilon}$
@@
$\min(\bm{R}_{t}(x_{h,i}))\prec\min(\bm{R}_{t}(y^{*}))~.$ (6)
Since $y^{*}\in{\cal S}_{t,0}\cup{\cal P}_{t,0}$, by Definition 8 and equation 6, we conclude that $x_{h,i}\notin{\cal P}_{pess,t}$. Thus, we must have $x_{h,i}\in S_{t,0}\setminus{\cal P}_{pess,t}$. Finally, we claim that $\max(\bm{R}_{t}(x_{h,i}))\preceq_{\bm{\epsilon}}\min(\bm{R}_{t}(y^{**}))$ for some $y^{**}\in{\cal P}_{pess,t}$. Since equation 5 implies that the condition $\max(\bm{R}_{t}(x_{h,i}))\preceq_{\bm{\epsilon}}\min(\bm{R}_{t}(y^{*}))$ is satisfied, then, if $y^{*}\in{\cal P}_{pess,t}$, we can simply set $y^{**}=y^{*}$. Else if $y^{*}\notin{\cal P}_{pess,t}$, then this will imply by Definition 8 existence of $y^{**}\in{\cal P}_{pess,t}$ such that $\min(\bm{R}_{t}(y^{*}))\prec\min(\bm{R}_{t}(y^{**}))$, which in turn together with equation 5 implies that $\max(\bm{R}_{t}(x_{h,i}))\prec_{\bm{\epsilon}}\min(\bm{R}_{t}(y^{**}))$. Then, by the discarding rule of Adaptive $\bm{\epsilon}$-PAL specified in line 9 of Algorithm 1, we must have $x_{h,i}$ discarded by the end of the discarding phase, which implies that $x_{h,i}$ cannot be in $S_{t,1}$. This proves that all $x_{h,i}\in{\cal S}_{t,0}\setminus{\cal P}_{t,2}$ must be discarded.∎
@@
B.3 Guarantees on the termination of Adaptive $\bm{\epsilon}$-PAL
We start by stating a bound on the maximum depth node that can be created by Adaptive $\bm{\epsilon}$-PAL before it terminates for a given $\bm{\epsilon}$. This result will be used in defining “good" events that hold with high probability under which $\bm{f}(x_{h,i})$ lies in the confidence hyper-rectangle of $x_{h,i}$ formed by Adaptive $\bm{\epsilon}$-PAL, for all possible nodes that can be created by the algorithm (see Section B.4). Our bounds on sample complexity will hold given that these “good" events happen.
@@ B.3 Guarantees on the termination of Adaptive $\bm{\epsilon}$-PAL
Proof. By Lemma 3, at the latest, the algorithm terminates at round $t$ for which $\overline{\omega}_{t}<\epsilon$, i.e., $\overline{\omega}^{2}_{t}<\epsilon^{2}$. By definition, at a refining round we have
@@
$16mV^{2}_{h_{t}}<\epsilon^{2}~,$
then we guarantee termination of the algorithm. Recall that we have defined
@@
for positive constants $C_{2}$ and $C_{3}$ defined in Corollary 1. Obviously, $V_{h}$ decays to 0 exponentially in $h$. Thus, by letting $h_{max}=h_{max}(\epsilon)$ to be the smallest $h\geq 0$, for which $16mV^{2}_{h_{max}}<\epsilon^{2}$ holds, it is observed that the algorithm stops refining at level $h_{max}$. ∎
@@
B.4 Two “good” events under which the sample complexity will be bounded
@@
(The first “good” event) For any $\delta\in(0,1)$, the probability of the following event is at least $1-\delta/2$:
@@
$1-\mathbb{P}(\mathcal{F}_{1})$ $=\mathbb{E}\left[\mathbb{I}\left(\exists j\in[m],\exists\tau\geq 0,\exists x\in\mathcal{T}_{h_{max}}:|f^{j}(x)-\mu^{j}_{\tau}(x)|>\beta^{1/2}_{\tau}\sigma^{j}_{\tau}(x)\right)\right]$
@@
(The second “good” event) For any $\delta\in(0,1)$, the probability of the following event is at least $1-\delta/2$:
@@
$1-\mathbb{P}(\mathcal{F}_{2})$ $=\mathbb{P}\left\{\exists h\geq 0,\exists i\in[N^{h}],\exists j\in[m]:\sup_{x\in B(x_{h,i},v_{1}\rho^{h},d)}|f^{j}(x)-f^{j}(x_{h,i})|>\omega(C_{\bm{k}}(v_{1}\rho^{h})^{\alpha})\right\}$
@@
B.5 Key results that hold under the “good” events
@@
In general, let the finite sequence of nodes $y_{1},y_{2},\ldots,y_{n}]^{\mathsf{T}}$ be such that all the nodes $y_{1}$ to $y_{n-1}$ are discarded, meaning that, for each $i\in[n-1]$, the node $y_{i}$ stopped being part of $\mathcal{P}_{pess,s_{i+1}}$ at some round $s_{i+1}$ by satisfying
$\min(\bm{R}_{s_{i+1}}(c_{s_{i+1}}(y_{i})))\prec\min(\bm{R}_{s_{i+1}}(y_{i+1}))~.$
@@
$\max(\bm{R}_{s_{1}}(c_{s_{1}}(x^{*})))-\bm{\epsilon}$ $\preceq\min(\bm{R}_{s_{1}}(y_{1}))\prec\min(\bm{R}_{s_{2}}(c_{s_{2}}(y_{1})))$
$\prec\min(\bm{R}_{s_{2}}(y_{2}))\prec\min(\bm{R}_{s_{3}}(c_{s_{3}}(y_{2})))$
$\prec\ldots$
$\prec\min(\bm{R}_{s_{n}}(y_{n}))\prec\min(\bm{R}_{t}(c_{t}(y_{n})))~,$
@@
$\max(\bm{R}_{s_{1}}(c_{s_{1}}(x^{*})))-\bm{\epsilon}\prec\min(\bm{R}_{t}(c_{t}(y_{n})))\preceq\max(\bm{R}_{t}(c_{t}(y_{n})))~.$
@@
Suppose that $z_{2}\notin\hat{P}$. Then, the above process continues in a similar fashion. Suppose that the process never yields a node in $\hat{P}$. Since the algorithm is guaranteed to terminate by Lemma 3, the process stops and yields a finite sequence $z_{1},\ldots,z_{n}]^{\mathsf{T}}$ of designs such that $z_{1}\notin\hat{P},\ldots,z_{n}\notin\hat{P}$. For each $i\in[n-1]$, let $s_{i+1}\in\mathbb{N}$ be the round at which $c_{s_{i+1}}(z_{i})$ is discarded; by the above process, we have
@@
$\min(\bm{R}_{s_{n+1}}(c_{s_{n+1}}(z_{n})))\prec\min(\bm{R}_{s_{n+1}}(z_{n+1}))~.$ (20)
At this point, it is important to clarify that $z_{n+1}$ cannot belong to the sequence $z_{1},\ldots,z_{n}]^{\mathsf{T}}$ of designs. Note that, by assumption, we have
$\bm{f}(x)$ $\preceq\min(\bm{R}_{s_{1}}(z_{1}))+\bm{\epsilon}\preceq\min(\bm{R}_{s_{2}}(c_{s_{2}}(z_{1})))+\bm{\epsilon}$
$\prec\min(\bm{R}_{s_{2}}(z_{2}))+\bm{\epsilon}\preceq\min(\bm{R}_{s_{3}}(c_{s_{3}}(z_{2})))+\bm{\epsilon}$
$\prec\ldots$
$\prec\min(\bm{R}_{s_{n}}(z_{n}))+\bm{\epsilon}\preceq\min(\bm{R}_{s_{n+1}}(c_{s_{n+1}}(z_{n})))+\bm{\epsilon}$
$\prec\min(\bm{R}_{s_{n+1}}(z_{n+1}))+\bm{\epsilon}~.$
Due to the strict domination relation between the nodes in the sequence (which is determined by Definition 8), no two nodes can have identical confidence hyper-rectangles at the time of removal from $\mathcal{P}_{pess,s}$, for $s\in\{s_{1},\ldots,s_{n+1}\}$. Moreover, such a condition implies a strict ordering of nodes in the above sense, which implies that $z_{n+1}\not\in\{z_{1},\ldots,z_{n}\}$. Hence, equation 20 and the earlier inequalities imply that
@@
where we have again used the shrinking property of the cumulative confidence hyper-rectangles in the second inequality. This means that $z_{n+1}$ is another node that $\bm{\epsilon}$-dominates $x$ but it is not in the finite sequence $z_{1},\ldots,z_{n}]^{\mathsf{T}}$. This contradicts our assumption that the process stops after finding $n$ designs. Hence, at least one of the designs in $z_{1},\ldots,z_{n}]^{\mathsf{T}}$ is in $\hat{P}$; let us call it $z$. This completes the proof of the claim.
@@
[…]
B.6 Derivation of the information-type sample complexity bound
@@
[…]
@@
$\sqrt{\frac{16\beta_{T}\sigma^{2}Cm\gamma_{T}}{T}}<\epsilon~.$
@@
Proof. According to Lemma 3, we have $\overline{w}_{t_{s}}<\epsilon$. In addition, Lemma 13 says that
@@
$T=\min\left\{\tau\in\mathbb{N}:\sqrt{\frac{16\beta_{\tau}\sigma^{2}Cm\gamma_{\tau}}{\tau}}<\epsilon\right\}~.$
Since the event $\{\tau_{s}=T\}$ implies that $\{\overline{w}_{t_{s}}<\epsilon\}$, we have $\Pr({\tau_{s}}>T)=0$.∎
B.7 Derivation of the dimension-type sample complexity bound